OpenStudy (anonymous):

Algebra 2: If p(x) = 2x2 – 4x and q(x) = x – 3, what is (pºq)(x)?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

The symbol in between p and q is the composite symbol. I do not know how to solve it.

OpenStudy (misty1212):

is that a little circle?

OpenStudy (misty1212):

kk we can do it

OpenStudy (misty1212):

\[(p\circ q)(x)=p(q(x))\]work from the inside out

OpenStudy (misty1212):

\[p(q(x))=p(x-3)\] is the first step to finding it

OpenStudy (anonymous):

Okay!

OpenStudy (misty1212):

then since \[p(\color\magenta\heartsuit)=2\color\magenta\heartsuit^2-4\color\magenta\heartsuit\] you have \[p(x-3)=2(x-3)^2-4(x-3)\]

OpenStudy (misty1212):

you can clean the last part up with some algebra

OpenStudy (anonymous):

Do you have to use distributive property??

OpenStudy (anonymous):

@misty1212